OpenStudy (loser66):

I need find z such that tan z = i/2 . I post my attempt and where I stuck. Please, help me out . BTW, the site is on and off at my end. Hence, if you don't see me, it doesn't mean I don't care your opinion. Thanks in advance.

OpenStudy (zarkon):

I get \[\frac{\ln(3)}{2}i\] solve \[\frac{e^{\theta}-e^{-\theta}}{e^{\theta}+e^{-\theta}}=\frac{1}{2}~\text{ ,for }~\theta\]

OpenStudy (loser66):

I got it. Thank you.

OpenStudy (loser66):

My solution \(tan z = \dfrac{sinz}{cos z}= \dfrac{\frac{e^{iz}-e^{-iz}}{2i}}{\dfrac{e^{iz}+e^{-iz}}{2}} =\dfrac{e^{iz}-e^{-iz}}{i(e^{iz}+e^{-iz})}=\dfrac{i}{2}\)

OpenStudy (loser66):

@freckles watch me, please

OpenStudy (loser66):

\(e^{iz}-e^{-iz} = -\dfrac{1}{2}(e^{iz}+e^{-iz})\)

OpenStudy (loser66):

\(2e^{iz}-2e^{-iz}= -e^{iz} -e^{-iz}\)

OpenStudy (loser66):

\(3e^{iz} -e^{-iz}=0\) multiple both sides by \(e^{iz}\) we have \(3e^{2iz} -1=0\\e^{2iz}= \dfrac{1}{3}\)

OpenStudy (freckles):

agree with everything you have so far

OpenStudy (loser66):

\(\implies 2iz = ln (1/3) \\z = \dfrac{ln(1/3)}{2i} = -i\dfrac{ln (1/3)}{2} = i\dfrac{ln 3}{2}\)

OpenStudy (loser66):

It is matching with what Zarkon gave me above

OpenStudy (loser66):

And next is my classmate solution.

OpenStudy (freckles):

should you write all the solutions... i*ln(3)/2-n*pi

OpenStudy (loser66):

how?

OpenStudy (freckles):

\[e^{2 i z}=e^{(2z)i}=e^{(2z+2 n \pi) i} \\ e^{(2 z + 2 n \pi )i}=\frac{1}{3} \\ (2 z + 2 n \pi) i =\ln(\frac{1}{3}) \\ 2 i (z + n \pi)=\ln(\frac{1}{3}) \\ 2i^2 (z+n \pi )=i \ln(\frac{1}{3} )\\ z+n \pi =\frac{-1}{2} i \ln (\frac{1}{3}) \\ z + n \pi =i \frac{1}{2}\ln((\frac{1}{3})^{(-1)}) \\ \\ z+n \pi =\frac{i \ln(3) }{2}\\ z=\frac{i \ln(3)}{2}- n \pi\]

OpenStudy (loser66):

Ok, now see what my classmate did. From \(e^{iz}-e^{-iz} = -\dfrac{1}{2}(e^{iz}+e^{-iz}) \) \(e^{iz}-e^{-iz} = -\dfrac{1}{2}(e^{iz})-\dfrac{1}{2}(e^{-iz})\) Let \(x = e^{iz}\) He got \(x - \dfrac{1}{x} = -\dfrac{1}{2}x - \dfrac{1}{2x}\)

OpenStudy (loser66):

That leads to the equation I asked and the solution is \(x =\pm\dfrac{\sqrt3}{3}\)

OpenStudy (loser66):

When plug back, he got \(e^{iz}= \pm\dfrac{\sqrt 3}{3}\) You see, it different from what we do above

OpenStudy (freckles):

\[e^{iz}= \frac{\sqrt{3}}{3} \\ e^{(z+2 n \pi)i}=\frac{\sqrt{3}}{3} \\ (z+2 n \pi)i=\ln(\frac{\sqrt{3}}{3}) \\ z+2n \pi =-i \ln(\frac{\sqrt{3}}{3}) \\ z+2 n \pi =i \ln( \frac{3}{\sqrt{3}}) \\ z+ 2 n \pi =i \ln(\sqrt{3}) \\ z=i \ln (\sqrt{3})-2 n \pi \\ \text{ and similarly solving } \\ e^{iz } =-\frac{\sqrt{3}}{3} \\ \text{ we get } \\ z=i \ln(-\sqrt{3})- 2 n \pi \\ \text{ but } \ln(-\sqrt{3})=\ln(3)+i \pi \\ \text{ so we have } \\ z=i \ln(\sqrt{3})+i \pi -2 n \pi \\ z=i \ln(\sqrt{3})-n \pi \]

OpenStudy (freckles):

the second solution there actually contains the first solution

OpenStudy (freckles):

one sec I made a mistake

OpenStudy (freckles):

\[\ln(-\sqrt{3})=\ln(\sqrt{3})+i \pi \text{ first I made a typeo here } \\ \text{ second I forgot \to do the following } \\ z=i \ln (-\sqrt{3})-2n \pi \\ z=i(\ln(\sqrt{3})+i \pi)-2n \pi \\ z=i \ln (\sqrt{3})- \pi -2n \pi \\ z=i \ln(\sqrt{3})-\pi (1+2n)\]

OpenStudy (freckles):

but anyways that solution gives us: \[z=i \ln(\sqrt{3})-\pi(1+2 n) \text{ or } z=i \ln(\sqrt{3})-\pi (2n)\] if I didn't do something wrong

OpenStudy (freckles):

and of course you can write i ln(sqrt(3)) as i/2 *ln(3)

OpenStudy (freckles):

that last thing I wrote is your solution and I believe the solutions are equivalent yours and your friend's

OpenStudy (loser66):

Thank you so much. We discuss but can't find what is wrong on both solutions.

OpenStudy (freckles):

is there suppose to be something wrong? is \[i \frac{1}{2} \ln(3)- n \pi \text{ not correct ?}\] or the other way to write it ... \[i \frac{1}{2} \ln(3)-2n \pi \text{ or } i \frac{1}{2}\ln(3)-(2n+1) \pi\]?

OpenStudy (freckles):

like I guess what I'm asking are you still seeing differences in the final solution ?

OpenStudy (loser66):

oh, I know how to Not redo my solution but still get Your answer From \(z = \dfrac{i}{2} ln 3\) we have \( |3|= 3, arg 3 =0 + 2k\pi\) hence \(ln 3 = ln 3 + i (2k\pi)\) Then \(z = \dfrac{i}{2} (ln 3 + i 2k\pi) = \dfrac{iln3}{2} - k\pi\)

OpenStudy (freckles):

ok cool stuff

OpenStudy (loser66):

:)

OpenStudy (loser66):

Again, thank you so much. Have a good night.

OpenStudy (freckles):

you too